### v2

# Tannakian categories: origins and summary

[…]
Tannakian categories: origins and summary
J.S. Milne
(February 23, 2025)
Origins
André Weil’s work on the arithmetic of curves and other varieties over finite fields led him in 1949 to state his famous “Weil conjectures”. These had a profound influence on algebraic geometry and number theory in the following decades. In an effort to explain the conjectures, Grothendieck was led to define several different “Weil cohomology theories” and to posit an ur-theory underlying all of them whose objects he called motives. In order to provide a framework for studying these different theories, especially motives, Grothendieck introduced the notion of a tannakian category.
Weil’s first insight was that the numbers of points on smooth projective algebraic varieties over finite fields behave as if they were the alternating sums of the traces of an operator acting on a well-behaved homology theory. [^1] Il me fallut du temps avant de pouvoir même imaginer que les nombres de Betti fussent susceptibles d’une interprétation en géométrie algébrique abstraite. Je crois que je fis un raisonnement heuristique basé sur la formule de Lefschetz. (It took me a while before I could even imagine that the Betti numbers were susceptible to an interpretation in abstract algebraic geometry. I think I made a heuristic argument based on the Lefschetz formula). Weil, Œuvre, Commentaire [1949b]. In particular, the (co)homology groups should be vector spaces over a field of characteristic zero, be functorial, and give the “correct” Betti numbers. However, already in the 1930s, Deuring and Hasse had shown that the endomorphism algebra of an elliptic curve over a field of characteristic $p$ may be a quaternion algebra over $\mathbb{Q}$ that remains a division algebra even when tensored with $\mathbb{Q}_{p}$ or $\mathbb{R}{}$, and hence cannot act on a $2$-dimensional vector space over $\mathbb{Q}{}$ (or even $\mathbb{Q}{}_{p}$ or $\mathbb{R}$). In particular, no such cohomology theory with $\mathbb{Q}{}$-coefficients exists.
Grothendieck defined étale cohomology groups with $\mathbb{Q}{}_{\ell}$-coefficients for each prime $\ell$ distinct from the characteristic of the ground field, and in characteristic $p\neq 0$, he defined the crystalline cohomology groups with coefficients in an extension of $\mathbb{Q}{}_{p}$. Each cohomology theory is well-behaved. In particular it has a Lefschetz trace formula, and Weil’s first insight is explained by realizing the points of the variety in a finite field as the fixed points of the Frobenius operator, and hence, by the trace formula, their cardinality as the alternating sum of the traces of the Frobenius operator acting on the cohomology groups. A striking feature of this is that, while the traces of the Frobenius operator are, by definition, elements of different fields $\mathbb{Q}{}_{l}$, they in fact lie in $\mathbb{Q}{}$ and are independent of $l$ (for smooth projective varieties). This last fact suggested to Grothendieck that there was some sort of $\mathbb{Q}{}$-theory underlying the different $\mathbb{Q}_{l}$-theories. To explain what this is, we need the notion of a tannakian category.
Briefly, a tannakian category over a field $k$ is a $k$-linear abelian category with a tensor product structure having most of the properties of the category of finite-dimensional representations of an affine group scheme over $k$ except one: there need not exist an exact tensor functor to the category of $k$-vector spaces, and when one does exist there need not be a canonical one. Each of the cohomology theories takes values, not just in a category of vector spaces, but in a tannakian category. For example, crystalline cohomology takes values in a category of isocrystals. These are finite-dimensional vector spaces over an extension of $\mathbb{Q}_{p}$, but only the elements of $\mathbb{Q}_{p}$ act as endomorphisms in the category. More specifically, if $\mathbb{1}$ is the unit object of the category (the tensor product of the empty set), we have $\mathrm{End}(\mathbb{1})=\mathbb{Q}_{p}$. Grothendieck’s insight is that there should be a tannakian category $\mathsf{Mot}$ over $\mathbb{Q}{}$ such that the functors to the local tannakian categories defined by the different cohomology theories factor through it. Algebraic correspondences between smooth projective algebraic varieties should define maps between motives, whose traces lie in $\mathrm{End}(\mathbb{1})=\mathbb{Q}$ and map to the traces on the various cohomology groups, which explains why the latter lie in $\mathbb{Q}$.
Weil’s second insight was that an analogue of the Riemann hypothesis should hold for the eigenvalues of Frobenius operators. This suggested that some of the well-known positivities in characteristic zero should persist to characteristic $p$. To see why, we briefly recall Weil’s proof of the Riemann hypothesis for abelian varieties over finite fields.
Consider an abelian variety $A$ over an algebraically closed field of characteristic $p$. For a prime $\ell\neq p$, we have a finite-dimensional $\mathbb{Q}{}_{\ell}$-vector space $V_{\ell}A$, and, for each polarization of $A$, we have a pairing $\varphi\colon V_{\ell}A\times V_{\ell}A\rightarrow\mathbb{Q}{}_{\ell}$. As $\mathbb{Q}{}_{\ell}$ is not a subfield of $\mathbb{R}{}$, it makes no sense to say that $\varphi$ is positive-definite. However, Weil showed that $\varphi$ induces an involution on the finite-dimensional $\mathbb{Q}{}$-algebra $\mathrm{End}(A)\otimes\mathbb{Q}{}$ and that this involution is positive. [^2] Over $\mathbb{C}$, this was known to the Italian geometers as the positivity of the Rosati involution. The Riemann hypothesis for the abelian variety follows directly from this. Grothendieck extended Weil’s ideas to tannakian categories by introducing the notion of a “Weil form” on an object of a tannakian category and of a “polarization” on a tannakian category.
A tannakian category over $k$ is said to be neutral if it admits an exact tensor functor to the category of $k$-vector spaces. Neutral tannakian categories are the analogues for affine group schemes of the categories studied by Tannaka and Krein. A classical theorem of Tannaka describes how to recover a compact topological group from its category of finite-dimensional unitary representations, and Krein characterized the categories arising in this way.
Not all tannakian categories are neutral, and the obstruction to a tannakian category over $k$ having a $k$-valued fibre functor lies in a nonabelian cohomology group of degree 2, more general than was available in the early 1960s. Grothendieck’s student Giraud developed the necessary nonabelian cohomology theory in his thesis. [^3] Giraud, Jean, Cohomologie non abélienne. Die Grundlehren der mathematischen Wissenschaften, Band 179. Springer-Verlag, Berlin-New York, 1971.
As we have explained, the idea of tannakian categories, and of their importance for motives, was Grothendieck’s. He explained it to Saavedra Rivano, who developed the theory of tannakian categories in his thesis. [^4] Saavedra Rivano, Neantro. Catégories Tannakiennes. Lecture Notes in Mathematics, Vol. 265. Springer-Verlag, Berlin-New York, 1972. It was Saavedra who introduced the terminology “tannakian”. Although Grothendieck used the term “tannakian category” in unpublished writings, he considered the categories to be part of a vast theory engobalizing Galois theory and the theory of fundamental groups, and later wrote that “Galois–Poincaré category” would have been a more appropriate name. [^5] Deligne writes: I expect that at first Grothendieck did not know of Tannaka’s work – and never cared about it. His aim was to unify the cohomology theories he had created. That each $H$ is with values in a category with $\otimes$, and that Künneth holds, was a brilliant insight which, like a number of his brilliant ideas, is now part of our subconscious, making it hard to see how deep it was.
Summary
We now present a summary of the main results of the theory. Throughout, $k$ is a field.
A tensor category (symmetric monoidal category) is a category $\mathsf{C}$ together with a functor $\otimes\colon\mathsf{C}{}\times\mathsf{C}\rightarrow\mathsf{C}$ and sufficient constraints to ensure that the tensor product of any (unordered) finite set of objects in $\mathsf{C}$ is well-defined up to a canonical isomorphism. In particular, there exists a unit object $\mathbb{1}$ (tensor product of the empty set of objects). A tensor category is rigid if every object admits a dual (in a strong sense). A tensor functor of tensor categories is one preserving the tensor products and constraints.
A tensorial category over $k$ is a rigid abelian tensor category equipped with a $k$-linear structure such that $\otimes$ is $k$-bilinear and the structure map $k\to\mathrm{End}(\mathbb{1})$ is an isomorphism. A tensorial category over $k$ is a tannakian category over $k$ if, for some nonzero $k$-algebra $R$, there exists an $R$-valued fibre functor, i.e., an exact $k$-linear tensor functor $\omega\colon\mathsf{C}\rightarrow\mathrm{Mod}(R)$. We write $\mathrm{Aut}^{\otimes}(\omega)$ for the group of automorphisms of $\omega$ (as a tensor functor).
In the remainder of the introduction, all tensor categories are assumed to be essentially small (i.e., equivalent to a small category).
A criterion to be a tannakian category
For an object $X$ of a tensorial category $\mathsf{C}$ over $k$, there is a canonical trace map
$\mathrm{Tr}_{X}\colon\mathrm{End}(X)\rightarrow\mathrm{End}(\mathbb{1})\simeq k,$
and we let $\dim X$ denote the trace of $\mathrm{id}_{X}$. In tensorial categories, traces are additive on short exact sequences (I, 6.6).  [^6] All references in this section are to: Milne, J. S., Tannakian Categories (version February 16, 2025), available here, from which this article has been adapted.
Theorem 1 (I, 10.1).
A tensorial category over $k$ of characteristic zero is tannakian (i.e., a fibre functor exists) if and only if, for all objects $X$, $\dim X$ is an integer $\geq 0$.
Neutral tannakian categories.
A tannakian category $(\mathsf{C},\otimes)$ over $k$ is neutral if there exists a $k$-valued fibre functor. For example, the category $\mathsf{Repf}(G)$ of finite-dimensional representations of an affine group scheme $G$ over $k$ is a tannakian category over $k$ with the forgetful functor as a $k$-valued fibre functor.
Theorem 2 (II, 3.1).
Let $\mathsf{C}$ be a tannakian category over $k$ and $\omega$ a $k$-valued fibre functor.
  1. (a)
    The functor of $k$-algebras $R\rightsquigarrow\mathrm{Aut}^{\otimes}(\omega\otimes R)$ is represented by an affine group scheme $G=\mathcal{A}ut^{\otimes}_{k}(\omega)$ over $k$.
  2. (b)
    The functor $\mathsf{C}\rightarrow\mathsf{Repf}(G)$ defined by $\omega$ is an equivalence of tensor categories.
For example, if $\mathsf{C}=\mathsf{Repf}(G)$ and $\omega$ is the forgetful functor, then $\mathcal{A}ut^{\otimes}_{k}(\omega)=G$.
The theorem gives a dictionary between neutralized tannakian categories over $k$ and affine group schemes over $k$. To complete the theory in the neutral case, it remains to describe the $R$-valued fibre functors on $\mathsf{C}$ for $R$ a $k$-algebra.
Theorem 3 (II, 8.1).
Let $\mathsf{C}$ and $\omega$ be as in Theorem 2, and let $G=\mathcal{A}ut^{\otimes}_{k}(\omega)$. For any $R$-valued fibre functor $\nu$ on $\mathsf{C}$, $\mathcal{I}som^{\otimes}(\omega\otimes R,\nu)$ is a torsor under $G_{R}$ for the fpqc topology. The functor $\nu\rightsquigarrow\mathcal{I}som^{\otimes}(\omega\otimes R,\nu)$ is an equivalence from the category of $R$-valued fibre functors on $\mathsf{C}$ to the category of $G_{R}$-torsors,
$\textsc{Fib}(\mathsf{C})_{R}\sim\textsc{Tors}(G)_{R}.$
Aside.
The situation described in the theorem is analogous to the following. Let $X$ be a connected topological space, and let $\mathsf{C}$ be the category of locally constant sheaves of $\mathbb{Q}{}$-vector spaces on $X$. For each $x\in X$, there is a functor $\omega_{x}\colon\mathsf{C}\rightarrow\mathsf{Vecf}_{\mathbb{Q}{}}$ taking a sheaf to its fibre at $x$, and $\omega_{x}$ defines an equivalence of categories $\mathsf{C}\rightarrow\mathsf{Rep}_{\mathbb{Q}}(\pi_{1}(X,x))$. Let $\Pi_{x,y}$ be the set of homotopy classes of paths from $x$ to $y$; then $\Pi_{x,y}\simeq\mathrm{Isom}(\omega_{x},\omega_{y})$, and $\Pi_{x,y}$ is a $\pi_{1}(X,x)$-torsor.
General tannakian categories.
Many of the tannakian categories arising in algebraic geometry are not neutral. They correspond to affine groupoid schemes rather than affine group schemes.
Let $S$ be an affine scheme over $k$. A $k$-groupoid scheme acting on $S$ is a $k$-scheme $G$ together with two $k$-morphisms $t,s\colon G\rightrightarrows S$ and a partial law of composition
[…]
such that, for all $k$-schemes $T$, $(S(T),G(T),(t,s),\circ)$ is a groupoid (i.e., a small category in which the morphisms are isomorphisms). The groupoid $G$ is transitive if the morphism
$(t,s)\colon G\rightarrow S\times_{k}S$
is faithfully flat. The representations of $G$ on locally free sheaves of finite rank on $S$ form a tannakian category $\mathsf{Repf}(S\mathrm{:}G)$ over $k$.
Let $S=\mathrm{Spec}(R)$ be an affine scheme over $k$. By a fibre functor over $S$, we mean an $R$-valued fibre functor. For example, $\mathsf{Repf}(S\mathrm{:}G)$ has a canonical (forgetful) fibre functor over $S$. When $\omega$ is a fibre functor over $S$ on a tannakian category over $k$, we let $\mathcal{A}ut_{k}^{\otimes}(\omega)$ denote the functor of $S\times_{k}S$-schemes sending $(b,a)\colon T\rightarrow S\times_{k}S$ to $\mathrm{Isom}_{T}^{\otimes}(a^{\ast}\omega,b^{\ast}\omega)$.
Theorem 4 (III, 1.1).
Let $\mathsf{C}$ be a tannakian category over $k$ and $\omega$ a fibre functor over $S$.
  1. (a)
    The functor $\mathcal{A}ut_{k}^{\otimes}(\omega)$ is represented by an affine $k$-groupoid scheme $G$ acting transitively on $S$.
  2. (b)
    The functor $\mathsf{C}\rightarrow\mathrm{Repf}(S\mathrm{:}G)$ defined by $\omega$ is an equivalence of tensor categories.
For example, if $\mathsf{C}=\mathrm{Repf}(S\mathrm{:}G)$ and $\omega$ is the forgetful functor, then $\mathcal{A}ut^{\otimes}_{k}(\omega)\simeq G$.
The gerbe of fibre functors
Let $\mathrm{Aff}_{k}$ denote the category of affine $k$-schemes. For each affine $k$-scheme $S$, we let $\textsc{Fib}(\mathsf{C})_{S}$ denote the category of fibre functors on $\mathsf{C}$ over $S$. As $S$ varies, the categories $\textsc{Fib}(\mathsf{C})_{S}$ form a stack over $\mathrm{Aff}_{k}$ for the fpqc topology, and (c) of Theorem 4 implies that $\textsc{Fib}(\mathsf{C})$ is a gerbe (any two fibre functors are locally isomorphic).
The tannakian categories over $k$ form a $2$-category with the $1$-morphisms being the exact $k$-linear tensor functors and the $2$-morphisms the morphisms of tensor functors. Similarly, the affine gerbes [^7] A gerbe is affine if the automorphisms of any object form an affine group scheme. over $k$ form a $2$-category with the $1$-morphisms being the cartesian functors of fibred categories and the $2$-morphisms being the equivalences between $1$-morphisms.
Theorem 5 (IV, 3.3).
The $2$-functor sending a tannakian category to its gerbe of fibre functors is an equivalence of $2$-categories. [^8] Not a $2$-equivalence Explicitly, for any tannakian category $\mathsf{C}$ over $k$, the canonical functor
$\mathsf{C}\rightarrow\mathrm{Repf}(\textsc{Fib}(\mathsf{C}))$
is an equivalence of tensor categories, and for any affine gerbe $\mathsf{G}$ over $k$, the canonical functor
$\mathsf{G}\rightarrow\textsc{Fib}(\mathrm{Repf}(\mathsf{G}))$
is an equivalence of stacks.
The theorem gives a dictionary between tannakian categories over $k$ and affine gerbes over $k$.
The fundamental group of a tannakian category
[…]
$\omega(\pi(\mathsf{T}))=\mathcal{A}ut^{\otimes}(\omega)$
for all fibre functors $\omega$. The group $\pi(\mathsf{T})$ acts on the objects $X$ of $\mathsf{T}$, and $\omega$ transforms this action into the natural action of $\mathcal{A}ut^{\otimes}(\omega)$ on $\omega(X)$.
Let $X$ be a topological space, connected, locally connected, and locally simply connected. There is the following analogy:
$\qquad\mathsf{T}$ $\qquad X$
object $Y$ of $\mathsf{T}$ covering of $X$(=locally constant sheaf)
fibre functor $\omega_{0}$ point $x_{0}\in X$
$\mathcal{A}ut^{\otimes}(\omega_{0})$ $\pi_{1}(X,x_{0})$
$\pi(\mathsf{T})$ local system of the $\pi_{1}(X,x)$
action of $\pi(\mathsf{T})$ on $Y$ in $\mathsf{T}$ action of the local system of the $\pi_{1}(X,x)$
on a locally constant sheaf.
For $\mathsf{T}$ the category of motives over $k$, $\pi(\mathsf{T})$ is called the motivic Galois group of $k$ [^9] From Deligne: The first three lines [in the table] were surely clear and important for Grothendieck. I don’t remember him considering $\mathrm{Ind}\mathsf{T}$, $\pi(\mathsf{T})$, or Hopf algebras in $\mathsf{T}$. For me, it was a way to make sense of my surprise, seeing that for each of the standard fibre functors $\omega$ with values in $\mathcal{C}$, $\mathcal{A}ut^{\otimes}(\omega\colon\text{motives}\to\mathcal{C}\to\text{vector spaces})$ had the same ‘texture’ as objects of $\mathcal{C}$.
Polarized tannakian categories.
For tannakian categories over $\mathbb{R}{}$ (or a subfield of $\mathbb{R}{}$), there are positivity structures called polarizations. For simplicity, let $(\mathsf{C},\otimes)$ be an algebraic tannakian category over $\mathbb{R}$. A nondegenerate bilinear form
$\phi\colon V\otimes V\rightarrow\mathbb{R}{}$
on an object $V$ of $\mathbb{C}{}$ is called a Weil form if its parity $\epsilon_{\phi}$ (the unique automorphism of $V$ satisfying $\phi(y,x)=\phi(x,\epsilon_{\phi}y)$) is in the centre of $\mathrm{End}(V)$ and if for all nonzero endomorphisms $u$ of $V$, $\mathrm{Tr}(u\circ u^{\phi})>0$, where $u^{\phi}$ is the adjoint of $u$. Two Weil forms $\phi\colon V\otimes V\rightarrow\mathbb{R}{}$ and $\psi\colon W\otimes W\rightarrow\mathbb{R}{}$ are compatible if the form $\phi\oplus\psi$ on $V\oplus W$ is again a Weil form.
Now fix an $\epsilon\in Z(\mathbb{R}{})$, where $Z$ is the centre of the band of the gerb of $\textsc{Fib}(\mathsf{C})$ – it is a commutative algebraic $\mathbb{R}{}$-group – and suppose that for each object $V$ of $\mathsf{C}$ we are given a nonempty compatibility class $\pi(V)$ of ($\pi$-positive) Weil forms on $V$ with parity $\epsilon_{V}$. We say that $\pi$ is an $\epsilon$-polarization of $\mathsf{C}$ if direct sums and tensor products of $\pi$-positive forms are $\pi$-positive. When $\epsilon=1$, so that $\phi(x,y)=\phi(y,x)$, the polarization is said to be symmetric.
Let $G$ be an affine group scheme over $\mathbb{R}$, and let $C$ be an element of $G(\mathbb{R})$ such that $\mathrm{inn}(C)$ is a Cartan involution, i.e., the involution corresponding to a compact form [^10] A real form $G^{\prime}$ of $G$ is compact if $G(\mathbb{R})$ is compact and contains a point of each connected component of $G_{\mathbb{C}}$. of $G$. Because $\mathrm{inn}(C)$ is an involution, $C^{2}$ is central. For each $V$ in $\mathsf{Repf}(G)$, let $\pi_{C}(V)$ be the set of $G$-invariant bilinear forms $\phi\colon V\otimes V\rightarrow\mathbb{R}{}$ such that the bilinear form $\phi_{C}$,
[…]
is symmetric and positive-definite. Then $\pi_{C}$ is a $C^{2}$-polarization on $\mathsf{Repf}(G)$. For a neutralized tannakian category, the $\pi_{C}$ exhaust the polarizations.
Theorem 6 (V, 8.2).
Let $G$ be an affine algebraic $\mathbb{R}{}$-group. Then $\mathsf{Repf}(G)$ admits a polarization if and only if $G$ is an inner form of a real compact group, in which case every polarization is of the form $\pi_{C}$ for some $C$ as above, and $C$ is uniquely determined by the polarization up to conjugacy.
It follows from the theorem that if $\mathsf{C}$ is an algebraic tannakian category endowed with a symmetric polarization, then $\mathsf{C}$ is neutral and there is an $\mathbb{R}{}$-valued fibre functor $\omega\colon\mathsf{C}{}\rightarrow\mathsf{Vecf}(\mathbb{R}{})$ such that $\mathcal{A}ut^{\otimes}(\omega)$ is a compact $\mathbb{R}$-group; moreover, $\omega$ is uniqely determined up to a unique isomorphism by the condition that the positive forms on an object $V$ of $\mathsf{C}$ are exactly the forms $\phi$ such that $\omega(\phi)$ is symmetric and positive-definite.
Motives
Fix an admissible equivalence relation for algebraic cycles on smooth projective algebraic varieties over $k$, and let $\mathsf{M}(k)$ denote the corresponding category of motives. It is a tensor category equipped with a $\mathbb{Q}$-linear structure (in particular, it is additive) such that $\otimes$ is $\mathbb{Q}$-bilinear.
Theorem 7 (VI, 2.5).
The category of motives $\mathsf{M}(k)$ is a $\mathbb{Q}$-linear rigid tensor category.
Let $X$ be a smooth projective variety over $k$. We say that $X$ satisfies the sign conjecture if there exists an algebraic cycle $e$ on $X\times X$ such that $e^{2}=e$ and $eH^{\ast}(X)=\bigoplus_{i\geq 0}H^{2i}(X)$ for the standard Weil cohomology theories. Smooth projective varieties over a finite field satisfy the sign conjecture, as do abelian varieties over any field. Let $\mathsf{NMot}(k)$ denote the category of motives for numerical equivalence over $k$ generated by the smooth projective varieties over $k$ satisfying the sign conjecture.
Theorem 8 (VI, 6.12).
The category of numerical motives $\mathsf{NMot}(k)$ is a semisimple tannakian category over $\mathbb{Q}$.
To prove that $\mathsf{NMot}(k)$ is polarized and that the standard Weil cohomologies factor through it requires Grothendieck’s standard conjectures. Given the lack of progress on these conjectures, Deligne has suggested looking for alternatives, of which there are several.
Acknowledgements
I thank Pierre Deligne for his generous help.